# Correlated Proxies: A New Definition and Improved Mitigation for Reward Hacking

### v2

@@ A.3 Action Distribution and Occupancy Measure Divergences in Contextual Bandits
That is, a single state is sampled from the initial state distribution $\mu_{0}$, and then a single action is sampled from the policy $\pi$ conditioned on that state. For example, RLHF for LLMs generally follows this setting: a prompt is sampled from a dataset, the LLM generates a response, and the reward is calculated based on the prompt and response.
@@
$D_{f}(\mu_{\pi}\;\|\;\mu_{\pi^{\prime}})=\mathbb{E}_{s\sim\mu_{0}(\cdot)}\left[D_{f}(\pi\;\|\;\pi^{\prime})\right].$
@@
A.4 State-only occupancy measures
@@ Appendix B Derivation of Occupancy-Regularized Policy Optimization
However, two terms in (11) cannot be computed directly: the current $\chi^{2}$ divergence between occupancy measures, and the ratio of the occupancy measures $\mu_{\pi_{\theta}}(s,a)/\mu_{\pi_{\text{base}}}(s,a)$. We first show how to approximate the latter using a discriminator network $\hat{d}^{\phi}(s,a)$, trained to optimize
@@
$\hat{d}^{\phi}(s,a)=\log\frac{\mu_{\pi_{\theta}}(s,a)}{\mu_{\pi_{\text{base}}}(s,a)}.$ (13)
@@
$\chi^{2}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{base}}}\right)=\mathbb{E}_{\mu_{\pi_{\theta}}}\left[\frac{\mu_{\pi_{\theta}}(s,a)}{\mu_{\pi_{\text{base}}}(s,a)}-1\right]=\mathbb{E}_{\mu_{\pi_{\theta}}}\left[e^{\hat{d}^{\phi}(s,a)}-1\right]\eqqcolon\widehat{\chi^{2}}$. (14)
@@
5:  Calculate $\widehat{\chi^{2}}=\mathbb{E}_{\mathcal{D}_{\pi}}\left[e^{\hat{d}^{\phi}(s,a)}-1\right].$
@@ C.2 Glucose monitoring
The RL agent (bottom) controls the insulin administered to a simulated patient in order to maintain healthy glucose levels. The true reward is a standard measure of health risk for the patient, but the proxy reward is misaligned and prioritizes the monetary cost of insulin. As the safe baseline policy, we train a BC policy based on data generated by a PID controller with parameters tuned by the original designers of the simulator (Steil, 2013).
@@ C.3 Pandemic mitigation
PandemicSimulator simulates a population’s infection dynamics using the SEIR model (Mwalili et al., 2020; Kompella et al., 2020):
@@ C.4 RLHF
We base our RLHF environment on the work of Coste et al. (2024), who study overoptimization of LLM-based reward models. The proxy reward model we use is fine-tuned from Pythia-70M (Biderman et al., 2023) on the AlpacaFarm (Dubois et al., 2023) preference dataset. As a true reward model, we use the Llama 3 Tulu V2 8B RM from AI2 (Ivison et al., 2024). As a base policy, we use Coste et al.’s SFT policy, which was fine-tuned from Python-1.4B on the AlpacaFarm SFT data. We evaluate policies’ true and proxy rewards on responses sampled for a held out set of prompts.
@@ Appendix E Additional Experiments and Results
E.1 Results for all regularization coefficients
@@ E.2 Ablations
Default parameters -1.15 $\pm$ 0.06 -12.18 $\pm$ 0.46 -47.5 $\pm$ 0.6 9.17 $\pm$ 0.12
Train policy before discriminator -1.24 $\pm$ 0.07 -12.23 $\pm$ 2.71 -48.2 $\pm$ 0.7 8.97 $\pm$ 0.13
Discriminator reward clipping $\times 0.1$ -1.18 $\pm$ 0.14 -12.22 $\pm$ 0.21 -47.7 $\pm$ 0.8 9.21 $\pm$ 0.18
Discriminator reward clipping $\times 10$ -1.24 $\pm$ 0.03 -12.11 $\pm$ 0.29 -47.9 $\pm$ 0.7 9.20 $\pm$ 0.14
Table 9: Results of our ablations of ORPO. We report the median and standard deviation of the true reward across five random seeds for the four non-RLHF environments. The top row shows the results of state-action occupancy measure regularization with the optimal coefficient for $\chi^{2}$ divergence. The other rows show ablations with the same coefficient. We find that ORPO is mostly robust to different hyperparameters but that it is probably best to train the discriminator network before the policy network.

### v3

@@ A.3 Action Distribution and Occupancy Measure Divergences in LLMs
That is, a single state is sampled from the initial state distribution $\mu_{0}$, and then a single action is sampled from the policy $\pi$ conditioned on that state. RLHF for LLMs follows this setting as a prompt is sampled from a dataset, the LLM generates a response, and the reward is calculated based on the prompt and response.
@@
$D_{f}(\mu_{\pi}\;\|\;\mu_{\pi^{\prime}})=\mathbb{E}_{s\sim\mu_{0}(\cdot)}\left[D_{f}\Big(\pi(\cdot\mid s)\;\|\;\pi^{\prime}(\cdot\mid s)\Big)\right].$
@@
A.3.1 Autoregressive Environments
While most RLHF implementations use the contextual bandit formulation above for the purposes of KL regularization, one can also model training an LLM as a sequential problem where each token generated is a separate action. This formulation is no longer a contextual bandit, but we can show that the action distribution and occupancy measure KL divergences are still equivalent!
Lemma A.7.
Suppose that an environment satisfies the following conditions:
  • It is deterministic: $\mu_{0}(s_{0})=1$ for exactly one state $s_{0}$, and for all $s_{t},a_{t}\in\mathcal{S}\times\mathcal{A}$, $p(s_{t+1}\mid s_{t},a_{t})=1$ for exactly one state $s_{t+1}$.
  • Exactly one sequence of actions leads to each state: if following $a_{0},\ldots,a_{t-1}$ leads to $s$, then no other sequence of actions (of any length) can also lead to $s$.
Then, for any policies $\pi,\pi^{\prime}$, the action distribution and occupancy measure KL divergences between them are equal (removing the $(1-\gamma)$ prefactor on the action distribution divergence):
$D_{\text{KL}}(\mu_{\pi}\;\|\;\mu_{\pi^{\prime}})=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}D_{\text{KL}}(\pi(\cdot\mid s_{t})\;\|\;\pi^{\prime}(\cdot\mid s_{t}))\right].$
Lemma A.7 applies to LLMs since one can treat the “state” of the environment after $t$ timesteps as all the tokens generated so far $w_{0}w_{1}\dots w_{t-1}$, and the actions as the next token $w_{t}$, which is then appended to the state:
$a_{t}\sim\pi(a_{t}\mid s_{t})$ $=\pi(w_{t}\mid w_{0}w_{1}\dots w_{t-1}).$
$p(s_{t+1}\mid s_{t},a_{t})$ $=\mathbf{1}\{s_{t+1}=w_{0}w_{1}\dots w_{t-1}w_{t}\}\qquad\text{where}\quad s_{t}=w_{0}w_{1}\dots w_{t-1}\quad\text{and}\quad a_{t}=w_{t}.$
Thus, regardless of the formalism used to train an LLM via RLHF, the action distribution and occupancy measure KL are equivalent. However, the conditions of Lemma A.7 are unlikely to be met by many other MDPs. Many MDPs are stochastic, violating the first assumption. Even among deterministic MDPs, it is very uncommon that only a single action sequence can lead to each state.
Proof.
Given the assumptions about the environment, we can rewrite the log-occupancy measure of a state-action pair in terms of the sum of log action probabilties over the unique sequence of actions leading to that state. Suppose $a_{0},\ldots,a_{t-1}$ is the unique action sequence leading to $s$ and that this action sequence visits states $s_{0},\ldots,s_{t-1},s$. Then
$\log\mu_{\pi}(s,a)$ $=\log\left((1-\gamma)\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathds{1}\{s_{t}=s\wedge a_{t}=a\}\right]\right)$
$=\log\left((1-\gamma)\gamma^{t}\mathbb{P}_{\pi}(s_{t}=s\wedge a_{t}=a)\right)$
$=\log\left((1-\gamma)\gamma^{t}\prod_{i=0}^{t}\pi(a_{i}\mid s_{i})\right)$
$=\log(1-\gamma)+t\log\gamma+\sum_{i=0}^{t}\log\pi(a_{i}\mid s_{i}).$
Using this, we can rewrite the occupancy measure KL divergence as
$D_{\text{KL}}(\mu_{\pi}\;\|\;\mu_{\pi^{\prime}})$ $=\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\mu_{\pi}(s,a)\log\left(\frac{\mu_{\pi}(s,a)}{\mu_{\pi^{\prime}}(s,a)}\right)$
$=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}\;\sum_{a_{0},\ldots,a_{t}\in\mathcal{A}^{t+1}}\;\,\mathbb{P}_{\pi}(a_{0}\wedge\cdots\wedge a_{t})\sum_{i=0}^{t}\Big(\log\pi(a_{i}\mid s_{i})-\log\pi^{\prime}(a_{i}\mid s_{i})\Big)$
$=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}\;\sum_{a_{0},\ldots,a_{t}\in\mathcal{A}^{t+1}}\;\,\left(\prod_{j=0}^{t}\pi(a_{i}\mid s_{i})\right)\sum_{i=0}^{t}\Big(\log\pi(a_{i}\mid s_{i})-\log\pi^{\prime}(a_{i}\mid s_{i})\Big),$ (17)
where $s_{i}$ is the state reached by taking $a_{0},\ldots,a_{i-1}$.
We will now show inductively that
$\sum_{a_{0},\ldots,a_{t}\in\mathcal{A}^{t+1}}\;\,\left(\prod_{j=0}^{t}\pi(a_{j}\mid s_{j})\right)\sum_{i=0}^{t}\Big(\log\pi(a_{i}\mid s_{i})-\log\pi^{\prime}(a_{i}\mid s_{i})\Big)$
$\qquad=\sum_{i=0}^{t}\sum_{s_{i}\in\mathcal{S}}\mathbb{P}_{\pi}(s_{i})D_{\text{KL}}(\pi(\cdot\mid s_{i})\;\|\;\pi^{\prime}(\cdot\mid s_{i})).$ (18)
Consider first if $t=0$. Then
$\sum_{a_{0}\in\mathcal{A}}\pi(a_{0}\mid s_{0})\Big(\log\pi(a_{0}\mid s_{0})-\log\pi^{\prime}(a_{0}\mid s_{0})\Big)$
$=D_{\text{KL}}(\pi(\cdot\mid s_{0})\;\|\;\pi^{\prime}(\cdot\mid s_{0}))$
$=\mathbb{P}_{\pi}(s_{0})D_{\text{KL}}(\pi(\cdot\mid s_{0})\;\|\;\pi^{\prime}(\cdot\mid s_{0})).$
Now suppose (A.3.1) holds for t1t-1. Then for $t$ we have
$\sum_{a_{0},\ldots,a_{t}\in\mathcal{A}^{t+1}}\;\,\left(\prod_{j=0}^{t}\pi(a_{j}\mid s_{j})\right)\sum_{i=0}^{t}\Big(\log\pi(a_{i}\mid s_{i})-\log\pi^{\prime}(a_{i}\mid s_{i})\Big)$
$=\sum_{a_{0},\ldots,a_{t-1}\in\mathcal{A}^{t}}\;\,\left(\prod_{j=0}^{t-1}\pi(a_{j}\mid s_{j})\right)\sum_{a_{t}\in\mathcal{A}}\pi(a_{t}\mid s_{t})\Bigg[\log\pi(a_{t}\mid s_{t})-\log\pi^{\prime}(a_{t}\mid s_{t})$
$\hskip 213.39566pt+\sum_{i=0}^{t-1}\Big(\log\pi(a_{i}\mid s_{i})-\log\pi^{\prime}(a_{i}\mid s_{i})\Big)\Bigg]$
$=\sum_{a_{0},\ldots,a_{t-1}\in\mathcal{A}^{t}}\;\,\left(\prod_{j=0}^{t-1}\pi(a_{j}\mid s_{j})\right)\Bigg[D_{\text{KL}}(\pi(\cdot\mid s_{t})\;\|\;\pi^{\prime}(\cdot\mid s_{t}))$
$\hskip 156.49014pt+\sum_{a_{t}\in\mathcal{A}}\pi(a_{t}\mid s_{t})\sum_{i=0}^{t-1}\Big(\log\pi(a_{i}\mid s_{i})-\log\pi^{\prime}(a_{i}\mid s_{i})\Big)\Bigg]$
$=\sum_{a_{0},\ldots,a_{t-1}\in\mathcal{A}^{t}}\;\,\left(\prod_{j=0}^{t-1}\pi(a_{j}\mid s_{j})\right)\Bigg[D_{\text{KL}}(\pi(\cdot\mid s_{t})\;\|\;\pi^{\prime}(\cdot\mid s_{t}))$
$\hskip 156.49014pt+\sum_{i=0}^{t-1}\Big(\log\pi(a_{i}\mid s_{i})-\log\pi^{\prime}(a_{i}\mid s_{i})\Big)\Bigg]$
$=\sum_{s_{t}\in\mathcal{S}}\mathbb{P}_{\pi}(s_{t})D_{\text{KL}}(\pi(\cdot\mid s_{t})\;\|\;\pi^{\prime}(\cdot\mid s_{t}))$
$\qquad+\sum_{a_{0},\ldots,a_{t-1}\in\mathcal{A}^{t}}\;\,\left(\prod_{j=0}^{t-1}\pi(a_{j}\mid s_{j})\right)\sum_{i=0}^{t-1}\Big(\log\pi(a_{i}\mid s_{i})-\log\pi^{\prime}(a_{i}\mid s_{i})\Big)$
$\overset{\text{(i)}}{=}\sum_{s_{t}\in\mathcal{S}}\mathbb{P}_{\pi}(s_{t})D_{\text{KL}}(\pi(\cdot\mid s_{t})\;\|\;\pi^{\prime}(\cdot\mid s_{t}))+\sum_{i=0}^{t-1}\sum_{s_{i}\in\mathcal{S}}\mathbb{P}_{\pi}(s_{i})D_{\text{KL}}(\pi(\cdot\mid s_{i})\;\|\;\pi^{\prime}(\cdot\mid s_{i}))$
$=\sum_{i=0}^{t}\sum_{s_{i}\in\mathcal{S}}\mathbb{P}_{\pi}(s_{i})D_{\text{KL}}(\pi(\cdot\mid s_{i})\;\|\;\pi^{\prime}(\cdot\mid s_{i})),$
where (i) is from the inductive hypothesis.
Now, plugging (A.3.1) into (17) gives
$D_{\text{KL}}(\mu_{\pi}\;\|\;\mu_{\pi^{\prime}})$
$=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}\sum_{i=0}^{t}\sum_{s_{i}\in\mathcal{S}}\mathbb{P}_{\pi}(s_{i})D_{\text{KL}}(\pi(\cdot\mid s_{i})\;\|\;\pi^{\prime}(\cdot\mid s_{i}))$
$=(1-\gamma)\sum_{i=0}^{\infty}\sum_{t=i}^{\infty}\gamma^{t}\sum_{s_{i}\in\mathcal{S}}\mathbb{P}_{\pi}(s_{i})D_{\text{KL}}(\pi(\cdot\mid s_{i})\;\|\;\pi^{\prime}(\cdot\mid s_{i}))$
$=(1-\gamma)\sum_{i=0}^{\infty}\sum_{s_{i}\in\mathcal{S}}\mathbb{P}_{\pi}(s_{i})D_{\text{KL}}(\pi(\cdot\mid s_{i})\;\|\;\pi^{\prime}(\cdot\mid s_{i}))\sum_{t=i}^{\infty}\gamma^{t}$
$=(1-\gamma)\mathbb{E}_{\pi}\left[\sum_{i=0}^{\infty}D_{\text{KL}}(\pi(\cdot\mid s_{i})\;\|\;\pi^{\prime}(\cdot\mid s_{i}))\sum_{t=i}^{\infty}\gamma^{t}\right]$
$=(1-\gamma)\mathbb{E}_{\pi}\left[\frac{\gamma^{i}}{1-\gamma}\sum_{i=0}^{\infty}D_{\text{KL}}(\pi(\cdot\mid s_{i})\;\|\;\pi^{\prime}(\cdot\mid s_{i}))\right]$
$=\mathbb{E}_{\pi}\left[\gamma^{i}\sum_{i=0}^{\infty}D_{\text{KL}}(\pi(\cdot\mid s_{i})\;\|\;\pi^{\prime}(\cdot\mid s_{i}))\right],$
which is the desired result. ∎
A.4 Learned Reward Functions are $r$-Correlated
Proxy reward functions are often learned from data like ratings or preference comparisons, including in the case of RLHF. Here, we show that a learned reward function with low mean-squared error—a common objective in supervised learning—is $r$-correlated with the true reward function.
Lemma A.8.
Let $R$ be the true reward function and $\tilde{R}$ be a learned reward function. Suppose that $\mathbb{E}_{\mu_{{\pi_{\text{base}}}}}\left[(R(s,a)-\tilde{R}(s,a))^{2}\right]\leq\epsilon\sigma_{R}^{2}$. Then, the learned reward function is an $r$-correlated proxy with $r\geq 1-\epsilon$.
The assumption in Lemma A.8 is that the mean-squared error over the occupancy measure of the base policy is small. This can be achieved, for example, by learning $\tilde{R}$ via least-squares regression over a training dataset of state-action pairs sampled from the base policy. Many results in learning theory show that this results in a small mean-squared error over the distribution the training data was sampled from, i.e., exactly the assumption in Lemma A.8 (Koltchinskii, 2006).
Proof.
Throughout the proof, all expectations, variances, and covariances are with respect to $\mu_{{\pi_{\text{base}}}}$. We can rewrite the assumption using the bias-variance decomposition as
$\mathbb{E}\left[(R(s,a)-\tilde{R}(s,a))^{2}\right]$
$\quad=\mathrm{Var}\left[R(s,a)-\tilde{R}(s,a)\right]+\left(\mathbb{E}\left[R(s,a)\right]-\mathbb{E}\left[\tilde{R}(s,a)\right]\right)^{2}$
$\quad=\mathrm{Var}\left[R(s,a)\right]+\mathrm{Var}\left[\tilde{R}(s,a)\right]-2\,\mathrm{Cov}\left[R(s,a),\tilde{R}(s,a)\right]+\left(\mathbb{E}\left[R(s,a)\right]-\mathbb{E}\left[\tilde{R}(s,a)\right]\right)^{2}$
$\quad=\sigma_{R}^{2}+\sigma_{\tilde{R}}^{2}-2\,\mathrm{Cov}\left[R(s,a),\tilde{R}(s,a)\right]+\left(\mathbb{E}\left[R(s,a)\right]-\mathbb{E}\left[\tilde{R}(s,a)\right]\right)^{2}$
$\quad\leq\epsilon\sigma_{R}^{2}.$
Note that $\left(\mathbb{E}\left[R(s,a)\right]-\mathbb{E}\left[\tilde{R}(s,a)\right]\right)^{2}>0$, so we can rewrite the inequality as
$2\mathrm{Cov}\left[R(s,a),\tilde{R}(s,a)\right]\geq(1-\epsilon)\sigma_{R}^{2}+\sigma_{\tilde{R}}^{2}\geq(1-\epsilon)\left(\sigma_{R}^{2}+\sigma_{\tilde{R}}^{2}\right).$
Dividing both sides by $2\sigma_{R}\sigma_{\tilde{R}}$ gives
$\frac{\mathrm{Cov}\left[R(s,a),\tilde{R}(s,a)\right]}{\sigma_{R}\sigma_{\tilde{R}}}$ $\geq\frac{1-\epsilon}{2}\frac{\sigma_{R}^{2}+\sigma_{\tilde{R}}^{2}}{\sigma_{R}\sigma_{\tilde{R}}}$.
By the AM-GM inequality, $\frac{\sigma_{R}^{2}+\sigma_{\tilde{R}}^{2}}{2}\geq\sigma_{R}\sigma_{\tilde{R}}$, so
$\frac{\mathrm{Cov}\left[R(s,a),\tilde{R}(s,a)\right]}{\sigma_{R}\sigma_{\tilde{R}}}\geq 1-\epsilon,$
which is the desired result. ∎
A.5 State-Only Occupancy Measures
@@ Appendix B Derivation of Occupancy-Regularized Policy Optimization
However, two terms in (19) cannot be computed directly: the current $\chi^{2}$ divergence between occupancy measures, and the ratio of the occupancy measures $\mu_{\pi_{\theta}}(s,a)/\mu_{\pi_{\text{base}}}(s,a)$. We first show how to approximate the latter using a discriminator network $\hat{d}_{\phi}(s,a)$, trained to optimize
@@
$\hat{d}_{\phi}(s,a)=\log\frac{\mu_{\pi_{\theta}}(s,a)}{\mu_{\pi_{\text{base}}}(s,a)}.$ (21)
@@
$\chi^{2}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{base}}}\right)=\mathbb{E}_{\mu_{\pi_{\theta}}}\left[\frac{\mu_{\pi_{\theta}}(s,a)}{\mu_{\pi_{\text{base}}}(s,a)}-1\right]=\mathbb{E}_{\mu_{\pi_{\theta}}}\left[e^{\hat{d}_{\phi}(s,a)}-1\right]\eqqcolon\widehat{\chi^{2}}$. (22)
@@
5:  Calculate $\widehat{\chi^{2}}=\mathbb{E}_{\mathcal{D}_{\pi}}\left[e^{\hat{d}_{\phi}(s,a)}-1\right].$
@@
How accurate is the discriminator-based approximation?.  To determine whether using the discriminator results in an accurate approximation of $\chi^{2}$ and $KL$ OM divergences, we plot the output of the discriminator in the glucose environment versus the theoretically correct value in Figure 6. The results suggest that the discriminator accurately approximates the log ratio of the occupancy measures, which in turn allows for accurate approximations of the OM divergences.
[…]
Figure 6: The top plot shows the state occupancy measures of the base policy $\pi_{\text{base}}$} and a policy $\pi$ optimized with ORPO in the glucose environment. The bottom plot shows the exact log ratio of the occupancy measures $\log\mu_{\pi}(s)/\mu_{\pi_{\text{base}}}(s)$ versus the discriminator output $\hat{d}_{\phi}(s)$, which attempts to approximate it. We find the discriminator output to be a good approximation of the log ratio.
@@ C.2 Glucose Monitoring
The RL agent (bottom) controls the insulin administered to a simulated patient in order to maintain healthy glucose levels. The true reward is a standard measure of health risk for the patient, but the proxy reward prioritizes the monetary cost of insulin. As the safe baseline policy, we train a BC policy based on data generated by a PID controller with parameters tuned by the original designers of the simulator (Steil, 2013).
In Figure 2, we adapt a diagram from Pauley et al. (2022) to represent this environment.
@@ C.3 Pandemic Mitigation
PandemicSimulator (Kompella et al., 2020) models a population’s infection dynamics using a modified SEIR model that they designed to represent the COVID-19 pandemic:
@@ C.4 RLHF
We base our RLHF environment on the work of Coste et al. (2024), who study overoptimization of LLM-based reward models. The proxy reward model we use is fine-tuned from Pythia-70M (Biderman et al., 2023) on the AlpacaFarm (Dubois et al., 2023) preference dataset. As a true reward model, we use the Llama 3 Tulu V2 8B RM from AI2 (Ivison et al., 2024). As a base policy, we use Coste et al.’s SFT policy, which was fine-tuned from Pythia-1.4B on the AlpacaFarm SFT data. We evaluate policies’ true and proxy rewards on responses sampled for a held out set of prompts.
@@ Appendix E Additional Experiments and Results
E.1 Win Rates for RLHF
Besides calculating the true reward for RLHF models with the Llama 3 Tulu V2 8B RM, we also calculate the win rates for the best coefficient of KL and $\chi^{2}$ regularization. We use AlpacaEval (Dubois et al., 2023) with GPT4o-mini to compute the win rate between the RLHF policy and the SFT policy. We find that the median win rate for $\chi^{2}$ divergence is higher, and it is also more consistent across random seeds. In contrast, using KL divergence leads to reward hacking for one seed and a lower median win rate.
Divergence Coefficient Median win rate Win rate range
$\chi^{2}$ 0.0008 52.83 51.98 – 53.98
KL 0.025 51.50 11.93 – 53.75
Table 4: Win rates for RLHF-trained models using KL divergence vs. $\chi^{2}$ divergence. The median win rate and range of win rates are reported across five seeds.
E.2 Results for All Regularization Coefficients
@@ E.3 Ablations
Default parameters -1.15 $\pm$ 0.06 -10.68 $\pm$ 0.15. -47.5 $\pm$ 0.6 9.17 $\pm$ 0.12
Train policy before discriminator -1.24 $\pm$ 0.07 -11.51 $\pm$ 0.40 -48.2 $\pm$ 0.7 8.97 $\pm$ 0.13
Discriminator reward clipping $\times 0.1$ -1.18 $\pm$ 0.14 -10.77 $\pm$ 0.16 -47.7 $\pm$ 0.8 9.21 $\pm$ 0.18
Discriminator reward clipping $\times 10$ -1.24 $\pm$ 0.03 -10.73 $\pm$ 0.37 -47.9 $\pm$ 0.7 9.20 $\pm$ 0.14
Table 10: Results of our ablations of ORPO. We report the median and standard deviation of the true reward across five random seeds for the four non-RLHF environments. The top row shows results using each environment’s most effective form of occupancy measure regularization: state-action OM for traffic, glucose, and tomato and state OM for pandemic, with the optimal $\chi^{2}$ divergence coefficients. The other rows show ablations with the same coefficient. We find that ORPO is mostly robust to different hyperparameters but that it is probably best to train the discriminator network before the policy network.